OpenStudy (anonymous):

Arlene wishes to invest 5000. if she invests part at 7% simple interest, part at 6% and receives a total of $332 after one year, how much does she invest at each rate?

OpenStudy (amistre64):

you just copying questions that already been answered?

OpenStudy (amistre64):

amount@7% = (332 - 5000(.06))/(.07-.06) amount@6% = (332 - 5000(.07))/(.06-.07)

OpenStudy (anonymous):

Sorry, OpenStudy made it double